# CLIP-QDA: An Explainable Concept Bottleneck Model

### v2

@@ 4.6 Assessing the Sample-wise Interpretability of XAI Methods
The objective of this subsection is to present our method within the context of current XAI methods applied to CBMs. We aim to compare CLIP-QDAglobal, CLIP-QDAlocal, CLIP-LIME and CLIP-SHAP with existing approaches such as SHAP (Lundberg & Lee, 2017), LIME (Ribeiro et al., 2016), LaBo (Yang et al., 2023), and Yan et al. (Yan et al., 2023b). A comparative overview of the distinctive features of these methods is presented in Table 3. Since Greybox XAI, X-NeSyL and Morales Rodríguez et al. require additional annotations, we chose to omit these methods from our study.
Table 3: Comparison of the different features of the XAI methods. Top: existing methods. Bottom: ours. Dataset-wise refers to methods that provide dataset-wise explanations. CLIP-QDAlocal refers to methods that provide sample-wise explanations. Closed-form solution refers to methods that do not require an optimization process to produce explanations. Computable from weights refers to methods that produce explanations from the model parameters and input values. Image level refers to methods that produce explanations from the image input. Concept level refers to methods that produce explanations from the concept input. Compatible with CLIP-QDA refers to methods that can produce explanations with CLIP-QDA as a classifier. Compatible with CLIP linear probe refers to methods that can produce explanations with CLIP linear probe as a classifier.
@@
To facilitate the comparison of various explanations generated for the same sample in the dataset, we computed the results for one image from PASCAL-Part (Donadello & Serafini, 2016). The results are categorized into three sections: sample-wise explanations at the image level, sample-wise explanations at the concept level, and dataset-wise explanations at the concept level. The subsequent section focuses on explaining the model’s prediction process given the image 12(a) as the input, which is labeled as “person”. The top two predicted labels by the classifier for this image are “person” and “potted plant”. Note that additional samples are available in Section A.7.
Image-level explanations are presented in Figure 12. For each of these samples, a weight is assigned to each pixel. All the results are computed by applying these post-hoc explanation methods on CLIP-QDA. Additional implementation and computation details can be found in A.1. While the results are easy to comprehend, criticism may arise due to potential misunderstandings regarding the specific focus of the method. While these methods successfully highlight the object of interest, it is challenging to discern the exact pattern the method emphasizes. Another observation is that, compared to classical classifiers, the results on CLIP-based CBMs are more imprecise, highlighting large areas of interest.
@@
Concept-level sample explanations are calculated in Figure 13. On a sample-wise basis, CLIP-LIME and CLIP-SHAP are computed using CLIP-QDA. Given that the method is model-specific, the Yan et al. method is computed on a linear classifier.
@@
Figure 13: Sample-wise explanations (concept level) of the image 12(a). Except for CLIP-SHAP, only the top 5 concepts are displayed. Note that the classifier correctly labeled the image.
First, we notice that there is a high diversity of concepts displayed among the methods. Overall, there are concepts associated with the prediction label (“Posture or stance”, “Human”) and concepts that arise from the specificities of the image (“Leaves”). Some concepts seem irrelevant (“Landing Gear”, “Windshield”, “Wooly”), which may be explained by the fact that their scores are useful to the classification process despite their absence in the image. Plus, it is worth noting that the context provided by the counterfactual analysis helps to unveil insights into the reason behind the explanation. For example, when taking into account the concept “Leaves” seems inappropriate, CLIP-QDAlocal explanations suggest that adding that concept influences the prediction of the class “potted plant,” indicating that this concept is important because it reveals the decision process behind predicting a person rather than a plant.
In the realm of dataset-wise explanations (Figure 14), we present the explanations for the CLIP-QDAglobal method, computed on CLIP-QDA, the LaBo explanation, computed on the LaBo classifier, and the Yan et al. explanation, computed on a linear classifier.
Figure 14: Dataset-wise explanations of the image 12(a). Only the top 5 concepts are displayed. Note that the classifier correctly labeled the image.
We observe that the concepts obtained from all the dataset-wise methods are generally coherent, including concepts commonly associated with persons, such as “Posture or stance” or “Facial features”. However, in contrast to the two other methods that impose positive magnitudes, the CLIP-QDAglobal explanation allows negative values, providing insights into whether the impact is negative or positive. Additionally, our CLIP-QDAglobal method emphasizes the comparison between two classes of interest, enabling a detailed exploration of specific disparities, like here between “potted plant” and “person”.
@@
To conduct a more comprehensive analysis beyond qualitative observations, we subject the various methods to quantitative evaluations. We apply the procedure outlined in Section 4.2 to the PASCAL-Part and Cats/Dogs/Cars datasets. In addition to CLIP-LIME, CLIP-SHAP, and QDA-CBM explanations, we also tested the explanations provided by LaBo and Yan et al. For the selection of $S_{s}$, in the case of the PASCAL-Part dataset, we opted to designate the concepts to detect as the concepts that are associated with the two highest probabilities of the model’s inference (refer to the Classes/Concepts association in Table 6). In the Cats/Dogs/Cars dataset, we chose to identify the concepts "Black" and "White" in the biased setup to mitigate potential biases. The resulting $Det$ score for this dataset is calculated as the average across all possible biased binary classification tasks. Additionally, we incorporated inference time as a parameter in our experiments referring to the time taken to produce explanations inferences on the entire validation set. Results on PASCAL-Part are displayed in Table 5 and results on Cats/Dogs/Car are displayed in Table 4.
@@
Regarding the Cats/Dogs/Cars dataset, our methods exhibit superior performance compared to state of the art, emphasizing the reliability of CLIP-QDA on datasets with a low number of concepts. Particularly, QDA-CBM demonstrates significantly faster inference times than CLIP-LIME and CLIP-SHAP, albeit at the expense of slightly lower deletion and detection scores. On datasets with a higher number of concepts and classes, such as PASCAL-Part, existing methods maintain higher scores.
@@ 5 Conclusion
In this paper, we introduce a modeling approach for the embedding space of CLIP, a foundation model trained on a vast dataset. Our observations reveal that CLIP can organize information in a distribution that exhibits similarities with a mixture of Gaussians. Building upon this insight, we develop an adapted concept bottleneck model that demonstrates competitive performance along with transparency. While the model that we have presented offers the advantage of simplicity and a limited number of parameters, it does encounter challenges when dealing with a broader and more ambiguous set of concepts. As a suggestion for future research, we propose to extend this modeling approach to incorporate other priors, such as Laplacian distributions, and to explore more complex models, including those with multiple components, i.e., using more than one Gaussian to describe a class for example. Another avenue to potentially enhance the performance of our model is to explore guiding the latent space towards a Gaussian distribution. Additionally, our research is centered around a specific embedding space (CLIP scores). Exploring similar work on other latent spaces, particularly those associated with multimodal foundation models, could be valuable to determine if similar patterns exist in those spaces.
@@ A.1 Implementation details
To compute GradCAM explanations, we applied the method to the $21^{th}$ block of the image encoder in CLIP-QDA.
@@
The LIME image-level explanations are generated using the superpixel approach defined in (Ribeiro et al., 2016). This method involves treating the input as a set of superpixels obtained through a segmentation process. LIME is then applied, considering these superpixels as binary values corresponding to whether they contribute to displaying the related zone. Following the original implementation, we use quickshift clustering to segment the image. We display the top 10 superpixels that exert the most influence on the prediction, highlighting superpixels in blue that have a negative impact on the label and superpixels in red that have a positive impact on the label.
@@
Following the advised method from the SHAP library, the image-level explanation from SHAP is computed using the partition explainer, which creates from the input a subset of images by applying a set of perturbations. In all our experiments, the perturbations are a blur of kernel (128,128).
@@
For the CLIP-LIME, we consider inputs as a set of tabular values. Then, each CLIP-LIME explanation is obtained by running the lime tabular explainer provided by the author’s repo with $num~{}features=100000$, $num~{}samples=1000$, and $batchsize=10$.
@@
For the CLIP-SHAP, we consider inputs as a set of tabular values. Then, each CLIP-SHAP explanation is obtained by running the shap kernel explainer provided by the author’s repo with a background of $5$ samples using k-means.
@@
For LaBo, we train a LaBo classifier by using Adam optimizer with a learning rate of $15e-$ 1 a weight decay of 00, and a batch size of $8192$. The resulting explanations follow the resulting weight matrix.
@@
For the Yan et al. method, we train a linear classifier by using Adam optimizer with a learning rate of $35e-$ 3 a weight decay of $41e-$ 4, and a batch size of $512$. The sample-wise (concept level) explanation results from the product between the concept score and its associated weight. The dataset-wise explanation results from the weights alone.
@@ A.2 Cats/Dogs/Cars
The Cats/Dogs/Cars dataset is available in https://huggingface.co/datasets/RemiKaz/Cats_Dogs_Cars. To create this dataset, we combined two well-known datasets: the Kaggle Cats and Dogs Dataset (Cukierski, 2013) and the Standford Cars Dataset (Krause et al., 2013). We then filtered the dataset to include only images featuring black and white animals and cars, leading to six different categories. The primary objective behind assembling this dataset is to facilitate research in scenarios characterized by significant data bias, such as the classification of white cats when the training data predominantly consists of images of white dogs and black cats.
@@ A.3 Set of concepts
Inspired by Yang et al. (2023), we use large language models to provide concept sets. Concretely, we use GPT-3 (Brown et al., 2020) with the following preprompt: “In this task, you have to give visual descriptions that describe an image. Respond as a list. Each item being a word.” Then, we generate the sets of words by the following prompt: “What are [N] useful visual descriptors to distinguish a [class] in a photo?” By doing so, we generated 5 concepts per class, presented in Table 6. An additional ordering of concepts by subcategories is available in Table 7.
@@ A.4 Study of images according to their position in the latent space
Figure 15: Images of minimal distance for each classes of PascalPART
@@
Figure 16: Images of maximal distance for each classes of PascalPART
@@
Figure 17: Images of minimal distance for each classes of MonuMAI
(a) “Baroque”
(b) ‘Gothic”
(c) “Hispanic-Muslim”
(d) “Renaissance”
Figure 18: Images of maximal distance for each classes of MonuMAI
For PascalPART, the closest images tend to be predominantly close-up shots of the object in question. However, this occasionally presents challenges due to the inherent pixelation of the images, leading to somewhat counterintuitive results. Conversely, the furthest images are typically those where the object is distant. Another scenario of failure arises when there are two classes present within the image, which complicates the retrieval process. On the other hand, for MonumAI, the closest images appear to be the simplest ones, without sign of disturbance. In contrast, the farthest images may include extraneous objects, such as signs.

### v3

@@ 5.3 Assessing the Sample-wise Interpretability of XAI Methods
The objective of this subsection is to present our method within the context of current XAI methods (see Section 4.2) applied to CBMs. A comparative overview of the distinctive features of these methods is presented in Table 3. Since Greybox XAI, X-NeSyL and Morales Rodríguez et al. require additional annotations, we chose to omit these methods from our study.
Table 3: Comparison of the different features of the XAI methods. Top: existing methods. Bottom: ours. Dataset-wise refers to methods that provide dataset-wise explanations. Sample-wise refers to methods that provide sample-wise explanations. Closed-form solution refers to methods that do not require an optimization process to produce explanations. Computable from weights refers to methods that produce explanations from the model parameters and input values. Image level refers to methods that produce explanations from the image input. Concept level refers to methods that produce explanations from the concept input. Compatible with CLIP-QDA refers to methods that can produce explanations with CLIP-QDA as a classifier. Compatible with CLIP linear probe refers to methods that can produce explanations with CLIP linear probe as a classifier.
@@
To facilitate the comparison of various explanations generated for the same sample in the dataset, we computed the results for one image from PASCAL-Part (Donadello & Serafini, 2016). The results are categorized into two sections: image-level explanations and concept-level explanations. The subsequent section focuses on explaining the model’s prediction process given the image 10(a) as the input, which is labeled as “person”. The top two predicted labels by the classifier for this image are “person” and “potted plant”. Note that additional samples are available in Section A.7.
Image-level explanations are presented in Figure 10. For each of these samples, a weight is assigned to each pixel. All the results are computed by applying these post-hoc explanation methods on CLIP-QDA. Additional implementation and computation details can be found in A.1. While the results are easy to comprehend, criticism may arise due to potential misunderstandings regarding the specific focus of the method. While these methods successfully highlight the object of interest, it is challenging to discern the exact pattern that the method emphasizes. Another observation is that, compared to classical classifiers, the results on CLIP-based CBMs are more imprecise, highlighting large areas of interest.
@@
Concept-level sample explanations are computed in Figure 11. On a sample-wise basis, CLIP-LIME and CLIP-SHAP are computed using CLIP-QDA. Given that the method is model-specific, the Yan et al. method is computed on a linear classifier.
@@
Figure 11: Sample-wise explanations (concept level) of the image 10(a). Except for CLIP-SHAP, only the top 5 concepts are displayed. Note that the classifier correctly labeled the image.
First, we notice that there is a high diversity of concepts displayed among the methods. Overall, there are concepts associated with the prediction label (“Posture or stance”, “Human”) and concepts that arise from the specificities of the image (“Leaves”). Some concepts seem irrelevant (“Landing Gear”, “Windshield”, “Wooly”), which may be explained by the fact that their scores are useful to the classification process despite their absence in the image. In addition, it is worth noting that the context provided by the counterfactual analysis helps to unveil insights into the reasons behind the explanation. For example, when taking into account the concept “Leaves” seems inappropriate, CLIP-QDAlocal explanations suggest that adding that concept influences the prediction of the class “potted plant,” indicating that this concept is important because it reveals the decision process behind predicting a person rather than a plant.
@@
To conduct a more comprehensive analysis beyond qualitative observations, we subject the various methods to quantitative evaluations. We apply the procedure outlined in Section 4.3 to the PASCAL-Part and Cats/Dogs/Cars datasets. In addition to CLIP-LIME, CLIP-SHAP, and QDA-CBM explanations, we also tested the explanations provided by LaBo and Yan et al. For the selection of $S_{s}$ in Equation 8, in the case of the PASCAL-Part dataset, we opted to designate the concepts to detect as the concepts that are associated with the two highest probabilities of the model’s inference (refer to the Classes/Concepts association in Table 6). In the Cats/Dogs/Cars dataset, we chose to identify the concepts “Black” and “White” in the biased setup to spot potential biases. The resulting $Det$ score for this dataset is calculated by considering $S_{s}=\left[Black,White\right]$ for classifiers trained with biased setups as described in 1. To reduce uncertainties, we do not only take the $Det$ score for the set [Black Cats, White Cars] but also for all possible similar biased binary classification tasks ([Black Cats, White Cars], [Black Dogs, White Cars], etc.). Additionally, we incorporated inference time as a parameter in our experiments referring to the time taken to produce explanations inferences on the entire validation set. Results on PASCAL-Part are displayed in Table 5 and results on Cats/Dogs/Cars are displayed in Table 4.
@@
Regarding the Cats/Dogs/Cars dataset, our methods exhibit superior performance compared to the state of the art, emphasizing the reliability of CLIP-QDA on datasets with a low number of concepts. Particularly, QDA-CBM demonstrates significantly faster inference times than CLIP-LIME and CLIP-SHAP, albeit at the expense of slightly lower deletion and detection scores. On datasets with a higher number of concepts and classes, such as PASCAL-Part, existing methods maintain higher scores.
5.4 Assessing the Dataset-wise Interpretability of XAI Methods
5.4.1 Evaluation of CLIP-QDAglobal on the PascalPART dataset
Following the setup expressed in Section 5.3, for dataset-wise explanations (Figure 12), we present the explanations for the CLIP-QDAglobal method, computed on CLIP-QDA, the LaBo explanation, computed on the LaBo classifier, and the Yan et al. explanation, computed on a linear classifier. Additional samples are available in Section A.7
Figure 12: Dataset-wise explanations of the image 10(a). Only the top 5 concepts are displayed. Note that the classifier correctly labeled the image.
We observe that the concepts obtained from all the dataset-wise methods are generally coherent, including concepts commonly associated with persons, such as “Posture or stance” or “Facial features”. However, in contrast to the other two methods that impose positive magnitudes, the CLIP-QDAglobal explanation allows for negative values, providing insights into whether the impact is negative or positive. Additionally, our CLIP-QDAglobal method emphasizes the comparison between two classes of interest, enabling a detailed exploration of specific disparities, like here between “potted plant” and “person”.
5.4.2 Evaluation of CLIP-QDAglobal on the Cats/Dogs/Cars dataset
We also present a toy example from our Cats/Dogs/Cars dataset by constructing a CBM consisting of the concepts Table 6, plus the concepts “Black” and “White”. Next, we display the 10 most influential concepts according to our global metric (the top 10 concepts that have the highest Wasserstein distance) in both biased and unbiased setups. Results are presented in Figure 13.
(a) Unbiased setup
(b) Biased setup
Figure 13: Global explanation on subsets of Cats/Dogs/Cars. Here, $c_{1}$=“Cat” and $c_{2}$=“Car”. Positive values indicate concepts that are more prevalent in cat images than car images, while negative values indicate concepts that are more common in car images compared to cat images. We display here only the top 10 concepts that have the highest Wasserstein distance (the concept “Black” is positioned $15^{th}$ in the unbiased setup).
Finally, we show an application of our local metric within the framework of our biased setup, as previously described. Subsequently, we feed an image of a white cat into our classifier (Figure 14). It is noteworthy that the image is misclassified as a car. Our local metric demonstrates sensitivity to the dataset’s color bias, corroborating the warning issued by the global explanation.
(a) Input image
Figure 14: Local explanation on subsets of Cats/Dogs/Cars. On the right figure, the blue/red scale represents the scaled counterfactuals as in equation 6, the text in each box corresponds to the label predicted after the perturbation, followed to the concept changed to obtain its result (in parentheses).
@@ 6 Conclusion
In this paper, we introduce a modeling approach for the embedding space of CLIP, a foundation model trained on a vast dataset. Our observations reveal that CLIP can organize information in a distribution that exhibits similarities with a mixture of Gaussians. Building upon this insight, we develop an adapted concept bottleneck model that demonstrates competitive performance along with transparency. While the model that we have presented offers the advantage of simplicity and a limited number of parameters, it does encounter challenges when dealing with a broader and more ambiguous set of concepts. As a suggestion for future research, we propose to extend this modeling approach to incorporate other priors, such as Laplacian distributions, and to explore more complex models, including those with multiple components, i.e., using more than one Gaussian to describe a class for example. Another avenue to potentially enhance the performance of our model is to explore guiding the latent space of CLIP’s scores towards a Gaussian distribution. Additionally, our research is centered around a specific embedding space (CLIP scores). Exploring similar work on other latent spaces, particularly those associated with multimodal foundation models, could be valuable to determine if similar patterns exist in those spaces.
7 Acknoledgements
This work was performed using HPC resources from GENCI-IDRIS (Grant 2023 - AD011014675).
@@ A.1 Implementation details
To compute GradCAM explanations, we applied the method to the $21^{th}$ block of the image encoder in CLIP-QDA, using the PyTorch package pytorch-grad-cam (jacobgilpytorchcam). Specifically, we generated the heatmap by backpropagating the gradient from the classifier’s inference class, considering the entire network (CLIP+classifier).
@@
The LIME image-level explanations are generated using the superpixel approach defined in (Ribeiro et al., 2016) using the official author’s repository. All of our samples are generated using the default parameters for image-level explanations, i.e. an exponential kernel of width of 0.25 on an image segmented using quickshift clustering. Concerning the visualization, each explanation underlines the top 10 superpixels that exert the most influence on the prediction. Superpixels that positively contribute to the prediction of the label are in red and superpixels that positively contribute to the prediction of the label are in blue.
@@
The SHAP image-level explanations are generated using the partitioning approach defined in the official author’s repository. All of our samples are generated using the default parameters for image-level explanations, i.e., running a blur image masker to partition the input image with rectangular masks. The estimation of Shapley values is obtained by running 200 evaluations. Partitions that positively contribute to the prediction of the label are in red and superpixels that negatively contribute to the prediction of the label are in blue.
@@
The CLIP-LIME explanations are generated using the approach used for tabular data as defined in (Ribeiro et al., 2016) using the official author’s repository. All of our samples are generated using the default parameters for tabular-level explanations, i.e., an exponential kernel of width 0.75×$\times N$, ($N$ being the number of concepts).
@@
The CLIP-SHAP explanations are generated using the approach used for tabular data as in the official author’s repository. All of our samples are generated using the default parameters for explanations, i.e., DeepLIFT algorithm using the training data as a background dataset.
@@
For LaBo, we train a LaBo classifier by using Adam optimizer with a learning rate of $0.5$, a weight decay of 00, and a batch size of $8192$. The resulting explanations follow the resulting weight matrix.
@@
For the Yan et al. method, we train a linear classifier by using Adam optimizer with a learning rate of $5\times 10^{-3}$, a weight decay of $10^{-4}$, and a batch size of $512$. The sample-wise (concept level) explanation results from the product between the concept score and its associated weight. The dataset-wise explanation results from the weights alone.
Resnet.
To train the Resnet classifier on PASCAL-Part, MIT scenes and MonuMAI, we initialized the network with ImageNet pertaining. Then, we trained the network using Adam optimizer with a learning rate of $10^{-3}$ for the probe, $10^{-4}$ for the backbone, a batch size of $64$, a momentum of $0.9$ and a weight decay of $10^{-4}$. For ImageNet results, we use results provided by the authors.
ViT.
To train the ViT classifier on PASCAL-Part, MIT scenes and MonuMAI, we initialized the network with ImageNet pertaining. Then, we trained the network using Adam optimizer with a learning rate of $10^{-2}$ for the probe, a frozen backbone, a batch size of $128$, a momentum of $0.9$, and no weight decay. For ImageNet results, we use results provided by the authors.
@@ A.2 Cats/Dogs/Cars
The Cats/Dogs/Cars dataset is available on hugging face. To create this dataset, we combined two well-known datasets: the Kaggle Cats and Dogs Dataset (Cukierski, 2013) and the Standford Cars Dataset (Krause et al., 2013). We then filtered the dataset to include only images featuring black and white animals and cars, leading to six different categories. The primary objective behind assembling this dataset is to facilitate research in scenarios characterized by significant data bias, such as the classification of white cats when the training data predominantly consists of images of white dogs and black cats.
@@ A.3 Set of concepts
Inspired by Yang et al. (2023), we use large language models to provide concept sets. Concretely, we use GPT-3 (Brown et al., 2020) with the following preprompt: “In this task, you have to give visual descriptions that describe an image. Respond as a list. Each item being a word.” Then, we generate the sets of words by the following prompt: “What are [N] useful visual descriptors to distinguish a [class] in a photo?”. By doing so, we generated 5 concepts per class, presented in Table 6. An additional ordering of concepts by subcategories is available in Table 7.
@@ A.4 Study of images according to their position in the latent space
Figure 15: Images of minimal distance for each class of PascalPART. Each subfigure in the plot corresponds to the image of class $c$ from the test set that exhibits the minimal Mahalanobis distance, as defined in Equation 9, with respect to the Gaussian representation of the given class $c$.
@@
Figure 16: Images of maximal distance for each class of PascalPART. Each subfigure in the plot corresponds to the image of class $c$ from the test set that exhibits the maximal Mahalanobis distance, as defined in Equation 9, with respect to the Gaussian representation of the given class $c$.
(a) “Baroque”
(b) “Gothic”
(c) “Hispanic-Muslim”
(d) “Renaissance”
Figure 17: Images of minimal distance for each class of MonuMAI. Each subfigure in the plot corresponds to the image of class $c$ from the test set that exhibits the minimal Mahalanobis distance, as defined in Equation 9, with respect to the Gaussian representation of the given class $c$.
@@
Figure 18: Images of maximal distance for each class of MonuMAI. Each subfigure in the plot corresponds to the image of class $c$ from the test set that exhibits the minimal Mahalanobis distance, as defined in Equation 9, with respect to the Gaussian representation of the given class $c$.
For PascalPART, the closest images tend to be predominantly close-up shots of the object in question. However, this occasionally presents challenges due to the inherent pixelation of the images, leading to somewhat counterintuitive results. Conversely, the farthest images are typically those where the object is distant. Another scenario of failure arises when there are two classes present within the image, which complicates the retrieval process. On the other hand, for MonumAI, the closest images appear to be the simplest ones, without sign of disturbance. In contrast, the farthest images may include extraneous objects, such as signs.